Gravitational field in a region is as given as , E= 3x2y2 i + 2x3y j. Gravitational potential at point A (1,1) is 10 unit, then what will potential at point B (2,3)

Grade:12

1 Answers

Samyak Jain
4 years ago
We know that potentail difference between two points A and B is given by
VB – VA = – ∫E.dr with lower and upper limits as coordinates of points A and B respectively
and dr = dx i + dy j + dz k.
Now, E.dr = (3x2y2 i + 2x3j).(dx i + dy j + dz k) = 3x2y2 dx + 2x3y dy
                 = d(x3 y2)
So, VB – VA =  – ∫E.dr = – ∫d(x3 y2) = – [x3 y2](1,1)(2,3)
i.e.  VB – 10 = – (23.32 – 13.12)  =  – 71
\therefore  VB  =  – 71 + 10 = – 61.
Potential at point B is – 61 unit.